write the balance equation for the complete combustion, in oxygen of the gaseous compounds (i) an alkane and (ii) alkyne each of which contains n-carbon atoms in molecules. Equal volume of the two hydrocarbon , measured at 120 degree celcius ,were separetely completely combusted in oxygen. The volume steam produced in (i) and (ii) were in the ratio 5:2 . Determine the value of "n" and hence the formulae of the alkane and alkyne
1
Expert's answer
2020-07-29T06:32:13-0400

Balance equation of complete combusation, in oxygen of the

(i) alkane CnH2n+2 + (3n+1)/2 O2 = n CO2 + (n+1) H2O


(ii) alkyne CnH2n-2 + (3n-1)/2 O2 = n CO2 + (n-1) H2O



(n+1)/(n-1) = 5/2

3n = 7

n = 2.33 approximate 2 (Answer)

Formula of (i) alkane = C2H6 (ii) alkyne = C2H4